OpenStudy (anonymous):

What is the simplest form of the radical expression? 4 3^sqrt 3x + 5 3^sqrt 10x

OpenStudy (mosaic):

Can you use the Equation editor to rewrite the problem so it is clearer?

OpenStudy (anonymous):

4\[4 \sqrt[3]{3x} + 5\sqrt[3]{10x}\]

OpenStudy (anonymous):

ignore the 4 on the top

OpenStudy (anonymous):

\[4\sqrt[3]{3x}+5\sqrt[3]{10x}\]

OpenStudy (mosaic):

Is it the above or \(\Large 4(3)^{\sqrt{3x}} + 5(3)^{\sqrt{10x}}\) ?

OpenStudy (mosaic):

You can post a screenshot of the question if you'd like.

OpenStudy (anonymous):

No its exactly how i posted it rn

OpenStudy (mosaic):

\( \Large 4\sqrt[3]{3x}+5\sqrt[3]{10x} = 4\sqrt[3]{3}\sqrt[3]{x}+5\sqrt[3]{10}\sqrt[3]{x} = \sqrt[3]{x}\{4\sqrt[3]{3}+5\sqrt[3]{10}\} \)

OpenStudy (anonymous):

how do i simplify that?

OpenStudy (mosaic):

There is a missing curly braces, }, at the end above. IDK if they expect you to use the calculator and simplify the constant within the curly braces. Can't think of any further simplification. Do you have answer choices?

OpenStudy (anonymous):

Yes hold on :P

OpenStudy (anonymous):

\[9\sqrt[3]{13x}\]

OpenStudy (anonymous):

\[27\sqrt[3]{10x}\]

OpenStudy (anonymous):

and the other one just says d. not possible to simplify

OpenStudy (mosaic):

Choice c is missing.

OpenStudy (anonymous):

Choice C is wrong tho

OpenStudy (mosaic):

I'd go with d)

OpenStudy (anonymous):

okay thanks :)